The volume of a cube is given by V = s3. Find the length of a side of a cube if the Volume is 729 cm3.
Answer:
Show work in this space.

Answer by pradhyumna.kumar@gmail.com(62)

The volume of a cube is given by V = s3. Find the length of a side of a cube if the Volume is 729 cm3.

sol :: s^3 = 729
s^3 = 9^3
=> s = 9